the city airport had a total of 349 departures on Monday, Tuesday, and Wednesday, there were 29 more departures on Tuesday than on Monday, There were 16 fewer departures on Wednesday than Monday. How many departures occurred each day. Applied problem solving in algebra
Answer by josgarithmetic(39800)

m,t,w, Monday Tuesday Wednesday

m+t+w=349; t=m+29; w=m-16.

Three linear equations in three unknown variables. The formulas for t and w are both expressed
in terms of m. Use substitution into the first equation sum of total departures.